OpenStudy (anonymous):

I really dont get this - Given the function k(x) = x2, compare and contrast how the application of a constant, c, affects the graph. The application of the constant must be discussed in the following manners: k(x + c) k(x) + c k(cx) c • k(x)

OpenStudy (lena772):

me either. sorry.

OpenStudy (anonymous):

okay ?

OpenStudy (jdoe0001):

I gather you haven't covered transformations? hmmm well, you can check it at http://www.youtube.com/watch?v=3Q5Sy034fok let us change the k(x) for the sake of less ambiguity make it f(x) so f(x + c) f(x) + c f(cx) c • f(x) \(\large{ \textit{parent function }\implies {\color{red}{ x^2}}\\ \quad \\ \begin{array}{llll} f(x + c)\implies &({\color{red}{ x+c}})^2\\ \quad \\ &\qquad \uparrow \\ &\textit{horizonal shift, positive, thus to the left}\\ f(x) + c\implies {\color{red}{ x^2}}&+c\\ \quad \\ &\uparrow \\ &\textit{vertical shift, positive, thus upwards}\\ f(cx)\implies &({\color{red}{ c\cdot x}})^2\implies c^2x^2\\ \quad \\ &\uparrow \\ &\textit{shrinks the graph by a factor of }c^2\\ c \cdot f(x)\implies &c\cdot {\color{red}{ x^2}}\\ &\uparrow \\ &\textit{shrinks the graph by a factor of }c\\ \end{array}}\)

OpenStudy (jdoe0001):